# SGLDBench: A Benchmark Suite for Stress-Guided Lightweight 3D Designs

### v1

@@ 3.4 Infill Voxelization
While topology optimization and porous infill compute a binary material field on a 3D voxel grid, the other methods compute a 3D lattice structure composed of edges and nodes. To enable a meaningful comparison of the structural properties of all approaches, SGLDBench voxelizes these structures into a voxel grid. The grid resolution is selected automatically to represent edges with a minimum required voxels. SGLDBench computes for each edge the intersected voxels via the DDA line drawing algorithm [55]. These voxels are set to solid. Edges are thickened by setting for all these voxels their 26 neighboring voxels to solid.
@@
For all edge-based infill strategies, the material budget and the targeted edge thickness are prescribed. The methods are then conducted in a dichotomy manner to find the settings that matches closely the design specifications, i.e., the design process is run multiple times to find the design that matches the material consumption under the edge thickness constraint.
3.5 Layout Analysis and Comparison
Once a 3D material design is computed, SGLDBench uses the linear elasticity solver to compute the compliance of the design. Irrespectively of the used design strategy, this requires a final FEA to simulate the stress field.
The user can view the meshes and graph structures generated by Voronoi infill, PSL-guided infill, conforming lattice structure and volumetric Michell trusses with MATLAB’s mesh viewing operations. The material fields generated with density-based topology optimization and porous infill optimization can be visualized with MATLAB’s functionality to render isosurfaces in 3D scalar fields, providing immediate feedback of the geometry and thickness of the produced structures.
Over that, SGLDBench provides an advanced WebGL-based volume visualization module to perform isosurface rendering and a stress-based comparative analysis of designs. Iso-surfaces in a 3D material field are rendered via ray-casting, with screen-space ambient occlusion for improved perception of the 3D structure of a design. All surface visualizations in this paper have been produced with this option.
@@ 3.5 Layout Analysis and Comparison
In addition, SGLDBench’s visualization module conveys the differences between the principal stress directions in the solid design and the computed infill design. SGLDBench computes the stress field using the initial boundary conditions, and then generates an auxiliary grid in which each voxel stores a single deviation measure. Based on the ordering of the absolute values of the principal stresses in the initial solid and the infill, the measure indicates the deviation of the stress directions corresponding to the principal stresses with maximum absolute value. The resulting scalar field is visualized via direct volume rendering. SGLDBench uses a transfer function that maps directional deviations linearly to colors ranging from white (low deviation) over light brown (medium deviation) to dark brown (high deviation). Opacity is linearly increased from low to high deviations, yet it starts at a base of 0.3 to also make low deviation regions visible.
@@ 4 Experiments
We demonstrate the use of SGLDBench to generate and analyze lightweight designs for different models and boundary conditions. We use a model of a human femur (Bone), a machine part that is often seen in engineering applications (Part), and Cantilever, a popular benchmark model in topology optimization (see Figs.  2 and  11). All models are initially given as triangle meshes.
All experiments are performed on a desktop computer equipped with an Intel CPU (6-core Xeon W2235), 64 GB of RAM, and an NVIDIA RTX 2070 GPU with 8 GB of video memory. We select a mid-end architecture on purpose to demonstrate SGLDBench’s capabilities even on affordable hardware. The main memory limits the maximum number of simulation elements that can be used by SGLDBench to roughly 160 million elements when simulating a 3D stress field. For Cantilever at a voxel grid resolution of $860\times 430\times 430$ (159 million elements), SGLDBench solves the FEA linear system on the CPU in roughly 30 minutes, with 41 solver iterations until reaching the convergence threshold of $1.0\times 10^{-3}$.
Topology optimization and porous infill optimization require additional memory for updating the material distribution using numerical optimization. SGLDBench performs these optimizations with about 130 million simulation elements. This corresponds to a $800\times 400\times 400$ simulation grid with 386 million degrees of freedom. SGLDBench requires about 45 minutes and 67 minutes, respectively, per optimization iteration of topology optimization and porous infill optimization. The results in Fig. 11 are generated by topology optimization and porous infill optimization with respectively 30 and 280 optimization iterations. In comparison to an optimized OpenMP CPU implementation of topology optimization [37], SGLDBench is only about a factor of 1.6 slower when reproducing the same cantilever model with equal resolution of $640\times 320\times 320$ on a similar computing architecture, i.e., a 48-core Xeon architecture. This demonstrates the efficiency of the MATLAB computing environment in combination with SGLDBench’s element-wise computation structure.
Figure 11: Material fields generated with topology optimization (a) and porous infill optimization (b) for Cantilever using a voxel grid of resolution $800\times 400\times 400$. The left face of the cubic design domain is fixed, and a downward force is uniformly applied at the bottom-right edge.
@@
All designs for Bone in Fig. 2 are generated with a $384\times 256\times 512$ voxel grid, corresponding to 9.8 million simulation elements. Topology optimization and porous infill optimization, respectively, generate the infill in 1.5 hours and 15.1 hours. Porous infill requires a significantly higher number of optimization iterations and additional computations to enforce the local volume constraint. Due to the high geometric complexity of porous infills, generating these infill also requires a higher number of iterations for solving the linear FEA system in each optimization iteration.
SGLDBench needs about 10 minutes to generate the Voronoi infill and the PSL-guided infill. About half of the time is required to simulate the 3D stress field in the initial solid domain. With this stress field, executing the external codes to generate the conforming lattice infill and the MATLAB code to generate the volumetric Michell truss requires 5 minutes and 4.5 hours, respectively. Slightly less than one half of the execution time of the latter is used for smoothing the original stress field.
Figure 12: (a) Density-based topology optimization. (b) Porous infill optimization. (c) Voronoi infill. (d) PSL-guided material layout. (e) Conforming lattice structures. (f) Volumetric Michell trusses. All infills consume roughly the same volume fraction of 0.22.
@@ 4.1 Material Use as a Modelling Parameter
To compare the response of different design methods to changes in the consumed material, we repeat all experiments shown in Fig. 2 with reduced volume fraction. Instead of a volume fraction of 0.4, SGLDBench now uses a low volume fraction of roughly 0.2 to enforce finer and lesser support structures. For all approaches generating a graph structure, the thickness of the voxelized edges is kept constant. The results are shown in Fig. 12, indicating increasing sparseness of all designs, at decreasing stiffness and with vastly different topologies. Topology optimization and porous infill do not require any modification of their simulation parameters, yet the entire optimization needs to be rerun to generate the results. All other approaches can use the stress field in the initial solid domain, and need to rerun only those parts generating the graph structure from it.
In particular, with a lower material budget the PSL-guided infill shows surprisingly good relative performance, since it uses the material to generate support structures along the most dominant stress directions. While with the 2D motivational example the conforming lattice infill has higher stiffness than the Voronoi infill and the volumetric Michell trusses (see Fig.  1), with Bone the mechanical properties have reversed. One reason might be that the resolution of the conforming lattice is too low, causing many edges of the graph structure to be misaligned with the dominant stress directions. In addition to the alignment of edges with the major stress directions, adaptive porosity also makes an important contribution to reducing compliance. I.e., in regions of high stresses, more material should be accumulated to resist the strain. This might be one reason for the rather moderate stiffness of the Voronoi infill, which cannot accumulate more material in regions of high stress due to the low material budget. Furthermore, the Voronoi infill, per construction, does not align with the dominant stress directions.
Since the stress field in Bone is rather simple and does not contain many degeneracies, most features can be well maintained by the Michell trusses. This, however, changes with Cantilever, see Fig. 15. Since the smoothed stress field of Cantilever shows significant differences to the initial field, many infill edges deviate strongly from the initial stress directions. This results in lower stiffness compared to the other alternatives.
@@ 4.2 Infill Resolution and Compliance
To shed light on the relationship between compliance, volume fraction, and geometric infill resolution, SGLDBench is used to evaluate the stiffness of high-resolution lattice infills for Bone. The infills are computed with SGLDBench, but they can also be generated with any other approach that produces an infill as a tetrahedral/hexahedral grid or an edge-graph. In the latter case, the infill is uploaded to SGLDBench and voxelized into a voxel grid. The volume fraction, the edge thickness, and the voxel grid resolution are defined by the user. Finally, SGLDBench computes the infill’s compliance using one FEA iteration.
In Fig. 13, a voxelized conforming lattice infill and a Voronoi infill with $136,877$ and $396,458$ edges, respectively, are shown. SGLDBench generates the infills using the same boundary conditions as in Figs  2 and  12. SGLDBench voxelizes the infills into a $848\times 576\times 1200$ voxel grid and builds a finite element model from it. The voxelized versions of the infills are visualized using SGLDBench’s WebGL visualization module with isosurface ray-casting and ambient occlusions. For computing the compliance of the voxel models, about $380$ million degrees of freedom are solved. SGLDBench does this in roughly 45 minutes. Interestingly, with the same material budget, the stiffness doesn’t seem to depend on the geometric details of the infills, as indicated by Fig.  2.
Figure 13: High resolution conforming lattice infill (a) and Voronoi infill (b), using the same boundary conditions and material budget as in Fig. 2.
@@ 4.3 Infill Geometry
Given the different compliances of different types of infills for the same object and load condition, we use SGLDBench to shed light on the major causes of these differences. Therefore, we compare the interior structure of the graph-based infill approaches, as demonstrated with Bone in Fig. 14. A clip plane is used to reveal some interior parts of the PSL-guided infill, the conforming lattice structure and the volumetric Michell trusses.
@@
The PSL-guided infill serves as a reference that shows the major load transmission pathways. The edges of the conforming lattice structure align with the PSLs overall. In regions where the curvature of the PSLs changes significantly (see the inset in Fig. 14), however, the lattice loses its geometric regularity. The degeneracies in the stress field were the principal stress directions flip seem to indicate these irregularities. Volumetric Michell trusses, on the other hand, exhibit a very regular geometric structure, yet it shows less alignment with the original stress directions. This is because the original stress field is smoothed, which changes the stress field globally and removes in particular unwanted degeneracies.
Figure 14: From (a) to (c), via a clip plane the interior of the PSL-guided lattice infill, the conforming lattice structure and volumetric Michell trusses are revealed. The closeup view shows the PSLs in the selected region, with the major, medium, and minor PSLs, respectively, shown in ocher, green and blue.
@@
The visualization of the 3D conforming lattice structure indicates a number of stitching edges in the interior that do not align with any of the dominant stress directions. The number of longer sequences of edges that follow consistently one of the PSL trajectories is lower compared to the PSL-guided infill and the volumetric Michell trusses. The reason for this behavior might be that the 3D conforming lattice structure uses too much material to obtain a consistent edge-graph rather than accounting for stiffness. Oscillating load paths, i.e., edge sequences, might result in less effective load transitions and reduced stiffness.
@@ 4.4 Reproducing Stresses in the Solid Domain
With Part shown in Fig. 15, SGLDBench is used to analyze how well different designs reproduce the stress directions in the initial solid configuration. In this way, users can, for instance, analyze whether the transmission of forces through tension and compression is along the same load paths as in the initial solid. Since an infill design can, in principle, transmit forces along the most efficient load paths in the solid design or use different paths, revealing relationships between compliance and reproduction of stress directions can give further hints on how to improve infill designs.
@@
Fig. 16 shows visualizations of the auxiliary grids that are computed for each infill. Direct volume rendering with the color and opacity transfer functions described in Sec.  3.4 is used to emphasize regions with high deviation in brown. Topology optimization generates an infill that in many parts reproduces the major stress directions in the solid under load. However, also some regions with major deviations can be seen, especially some thin structures and structures close to the fixed elements where stresses are highest. The same holds for porous infill, yet overall it shows slightly higher deviations that are spread over the entire domain due to the more space-filling distribution of material.
It is clear that the stress deviations are highest in the Voronoi infill, because the initial stress directions are not considered in the edge-graph layout. The stress deviations are low when edges cover boundary elements, since these elements are considered in the stress analysis of the solid object and the infill. The material in the PSL-guided infill is distributed along the principal stress directions, and, thus, stresses are well aligned with these directions. However, as also seen in Fig. 3, at the boundaries of the voxelized edges the alignment is less accurate. This is due to inaccuracies in the stress simulation using the discretized line structures, a property that holds for all graph-based approaches when using the voxelized structures for the compliance analysis. Furthermore, since all graph-based infills contain many edges along the two stress directions not corresponding to the direction with the maximum absolute stress value, a high directional deviation can occur along these edges.
@@
The deviation maps of the conforming lattice structure and the Michell trusses are very similar, showing the same low deviation patterns in the boundary region as the Voronoi infill. In both infills there is still a considerable number of voxels with significantly different dominant stress direction than the solid object.
Figure 15: (a) Density-based topology optimization. (b) Porous infill optimization. (c) Voronoi infill. (d) PSL-guided infill. (e) Conforming lattice structures. (f) Volumetric Michell trusses. All infills consume roughly the same volume fraction of 0.3.
Figure 16: Stress-to-stress alignment. (a) Topology optimization. (b) Porous infill optimization. (c) Voronoi infill. (d) PSL-guided infill. (e) Conforming lattice. (f) Volumetric Michell trusses.
In the supplementary material, we further demonstrate the usability of SGLDBench with additional datasets, which will be made available together with the code repository.

### v2

@@ 3.4 Infill Voxelization
While TO and porous infill compute a binary material field on a 3D voxel grid, the other methods compute a 3D lattice structure composed of edges and nodes. To enable a meaningful comparison of the structural properties of all approaches, SGLDBench voxelizes these structures into a voxel grid. The grid resolution is selected automatically to represent edges with a minimum required voxels. SGLDBench computes for each edge the intersected voxels via the DDA line drawing algorithm [56]. These voxels are set to solid. Edges are thickened by setting for all these voxels their 26 neighboring voxels to solid.
@@
For all edge-based infill strategies, the material budget and the targeted edge thickness are prescribed. The methods are then conducted in a dichotomy manner to find the settings that match closely the design specifications, i.e., the design process is run multiple times to find the design that matches the material consumption under the edge thickness constraint.
3.5 Visualization and Layout Analysis
Once a Voronoi infill, PSL-guided infill, conforming lattice structures, or volumetric Michell trusses has been computed, users can view the meshes and graph structures using MATLAB’s mesh viewing operations, rendering edges as cylinders with a specified width. Thus, voxelizing the infill into a 3D material field is not required.
To compute the compliance of a design, lattice infill designs must first be voxelized into a 3D material field. SGLDBench then uses its linear elasticity solver to perform an FEA and simulate the stress field, from which the compliance is computed. For density-based TO and porous infill optimization, the compliance history throughout the optimization process is recorded and can be visualized via a curve plot, as demonstrated in Fig. 5.
To visualize a stress field, an evenly spaced set of PSLs covering the domain as uniformly as possible is computed using MATLAB programs. Users can control the density of seeded PSLs and select scalar stress measures, such as the von Mises norm, to map onto the lines’ colors. PSLs can be computed for the initial stress field in the solid domain. Examples of PSL-guided visualizations using MATLAB are shown in Fig. 8.
Figure 8: PSL visualization with cylindrical elements and color-coded von Mises stresses using MATLAB visualization programs.
For realtime visualization of even high resolution designs, SGLDBench provides an advanced WebGL-based volume visualization module. It performs isosurface and direct volume rendering, and assists users in a stress-based comparative design analysis. Isosurfaces in a 3D material field are rendered using GPU ray-casting, with screen-space ambient occlusion to enhance depth perception, as demonstrated in Fig. 1. A view-space parallel clip plane can be moved back and forth to expose otherwise occluded structures.
Direct volume rendering is particularly used to visualize the scalar von Mises stress field, enabling the evaluation of whether a material will permanently deform under the given stress state. High magnitudes of the von Mises stress indicate a risk of fracture under the applied loads, and direct volume rendering effectively highlights the spatial regions where this danger is significant. This provides a powerful tool for structural and mechanical analysis. SGLDBench’s WebGL interface allows users to color a ray-traced infill surface based on the von Mises stress.
@@ 3.5 Visualization and Layout Analysis
In addition, SGLDBench’s visualization module highlights the differences between the principal stress directions in the solid design and the computed infill design. SGLDBench computes the stress field of the infill using the initial boundary conditions and generates an auxiliary grid where each voxel stores a single deviation measure. This measure indicates the deviation of the stress directions corresponding to the principal stresses with the maximum absolute value, based on the ordering of the absolute values of the principal stresses in the initial solid and the infill.
The resulting scalar field is visualized through direct volume rendering. SGLDBench employs a transfer function that maps directional deviations linearly to colors, ranging from white (low deviation) to light brown (medium deviation) and dark brown (high deviation). Opacity is initially set to one but can be adjusted by the user to smoothly fade out regions with low or high deviation.
Given an optimized infill structure, SGLDBench can also be used to apply new boundary conditions to it, allowing users to probe conditions different from those for which the structure was initially optimized. This functionality is inspired by worst-case structural analysis [57, 58], a method used to evaluate the performance and reliability of a structure under its most unfavorable conditions. While SGLDBench is designed for a completely different use case and cannot perform such analysis, it provides the tools to explore similar scenarios.
Specifically, SGLDBench includes an interface to modify the initial boundary conditions by changing the direction of forces and re-computing the compliance and von Mises stress under the new conditions. A visualization of the structure with stress-based color coding highlights the mechanical strengths and weaknesses of different parts, enabling a deeper understanding of its behavior under varied conditions.
@@ 4 Experiments
We demonstrate the use of SGLDBench to generate and analyze lightweight designs for various models and boundary conditions. The models include a human femur (Bone), a machine part commonly seen in engineering applications (Part), and Cantilever, a widely used benchmark model in TO. All models are initially provided as triangle meshes. We showcase the usability of SGLDBench with additional datasets in the supplementary material.
All experiments are conducted on a desktop computer equipped with an Intel 6-core Xeon W2235 CPU, 64 GB of RAM, and an NVIDIA RTX 2070 GPU with 8 GB of video memory. We intentionally select a mid-range architecture to demonstrate SGLDBench’s capabilities on affordable hardware. The main memory limits the maximum number of simulation elements to approximately 160 million for simulating a 3D stress field.
4.1 Performance Evaluation
Solver. For the Cantilever model at a voxel grid resolution of 860×430×430 (159 million elements), SGLDBench solves the FEA linear system on the CPU in roughly 30 minutes, achieving convergence within 41 solver iterations at a threshold of $1.0\times 10^{-3}$
@@ 4.1 Performance Evaluation
Iterative Optimization. TO and porous infill optimization require additional memory for updating the material distribution during numerical optimization. SGLDBench performs these optimizations with grids of up to 130 million simulation elements, corresponding to a 800×400×400 simulation grid with 386 million degrees of freedom. For these cases, SGLDBench completes each optimization iteration in approximately 45 minutes for TO and 67 minutes for porous infill optimization. The results shown in Fig. 9 are generated using 30 iterations of TO and 280 iterations of porous infill optimization.
Compared to an optimized OpenMP CPU implementation of TO [38], SGLDBench is only about 1.6 times slower when reproducing the same Cantilever model at a resolution of 640×320×320 on a similar computing architecture (a 48-core Xeon CPU). This demonstrates the efficiency of the MATLAB computing environment in combination with SGLDBench’s element-wise computation structure.
Figure 9: Material fields generated with TO (a) and porous infill optimization (b) for Cantilever using a $800\times 400\times 400$ voxel grid. The left face of the cubic design domain is fixed, and a downward force acts along the bottom-right edge.
Figure 10: Same designs as in Fig. 1 but with lower material budget of roughly 0.22.
All designs for Bone in Fig. 1 are generated with a $384\times 256\times 512$ voxel grid, corresponding to 30 million degrees of freedom. TO and porous infill optimization, respectively, generate the infill in 1.5 hours and 15.1 hours. Porous infill requires a significantly higher number of optimization iterations and additional computations to enforce the local volume constraint. Due to the high geometric complexity of porous infills, generating these infill also requires a higher number of iterations for solving the linear FEA system in each optimization iteration.
Lattice Infill Optimization. To compute the various types of lattice infills, we utilize state-of-the-art implementations currently available, which are provided as either C/C++ codes (PSL-guided layouts, Voronoi, and conforming lattice infills) or MATLAB programs (Michell trusses). Executing these implementations through SGLDBench does not result in performance penalties.
SGLDBench requires approximately 10 minutes to generate the Voronoi infill and the PSL-guided infill, with about half of this time spent simulating the 3D stress field in the initial solid domain. Using this stress field, the external codes for generating the conforming lattice infill take around 5 minutes, while the MATLAB code for generating the volumetric Michell truss requires approximately 4.5 hours. For the latter, slightly less than half of the execution time is allocated to smoothing the original stress field.
@@ 4.2 Material Use as a Modelling Parameter
To evaluate the response of different design methods to changes in material consumption, we repeat all experiments shown in Fig. 1 with a reduced volume fraction. Instead of the original volume fraction of 0.4, SGLDBench now uses a lower volume fraction of approximately 0.2, enforcing finer and less dense support structures. For methods that generate graph structures, the thickness of the voxelized edges remains constant.
The results, shown in Fig. 10, reveal an increased sparseness in all designs, accompanied by reduced stiffness and significantly varied topologies. TO and porous infill methods require no changes to their simulation parameters but must rerun the entire optimization process to produce the results. In contrast, other approaches can utilize the stress field from the initial solid domain and need to rerun only the steps that generate the graph structure from it.
With a lower material budget, the PSL-guided infill demonstrates surprisingly good relative performance, as it effectively utilizes the material to generate support structures along the most dominant stress directions.
One possible reason for this reversal is that the resolution of the conforming lattice is too low, leading to the misalignment of many edges in the graph structure with the dominant stress directions. Furthermore, adaptive porosity also plays a critical role in reducing compliance. Specifically, in regions of high stress, more material should be allocated to resist strain effectively. This could explain the relatively moderate stiffness of the Voronoi infill, which is unable to concentrate more material in high-stress regions due to the low material budget. Moreover, by design, the Voronoi infill does not align with the dominant stress directions, further contributing to its reduced mechanical performance.
The stress field in Bone is relatively simple and contains few degeneracies, allowing the Michell trusses to preserve most features effectively. This behavior, however, changes with Part (see the first row of Fig. 13). In this case, the smoothed stress field exhibits significant differences from the initial field, causing many infill edges to deviate substantially from the original stress directions. As a result, the Michell trusses demonstrate lower stiffness compared to the other alternatives.
@@ 4.3 Infill Resolution and Compliance
To explore the relationship between compliance, volume fraction, and geometric infill resolution, SGLDBench is used to evaluate the stiffness of high-resolution lattice infills for Bone. The volume fraction, edge thickness, and voxel grid resolution are user-defined parameters, and SGLDBench computes the compliance of the infill using a single FEA iteration.
In Fig. 11, a voxelized conforming lattice infill and Voronoi infill with $136,877$ and $396,458$ edges, respectively, are shown, using the same boundary conditions as in Figs  1 and  10. SGLDBench generates a 848×576×1200 voxel grid and constructs a finite element model from it. The voxelized versions of the infills are visualized using SGLDBench’s WebGL visualization module, employing isosurface ray-casting and ambient occlusion for enhanced depth perception.
To compute the compliance of the voxel models, approximately 380 million degrees of freedom are solved, which SGLDBench completes in roughly 45 minutes. Interestingly, as shown in Fig. 1, the stiffness appears to be independent of the geometric details of the infills when using the same material budget.
Figure 11: High resolution results using the same boundary conditions and material budget as in Fig. 1.
@@ 4.4 Infill Geometry
Given the varying compliance of different types of infills for the same object and load conditions, we use SGLDBench to investigate the major causes of these differences. We compare the interior structures of graph-based infillS, as illustrated with Bone in Fig. 12. A clip plane is used to expose the interior regions of the PSL-guided infill, the conforming lattice structure, and the volumetric Michell trusses.
@@
The PSL-guided infill serves as a reference, illustrating the primary load transmission pathways. Overall, the edges of the conforming lattice structure align with the PSLs. However, in regions where the curvature of the PSLs changes significantly (see the inset in Fig. 12), the lattice loses its geometric regularity. These irregularities appear to be associated with degeneracies in the stress field, where the principal stress directions flip.
In contrast, the volumetric Michell trusses exhibit a highly regular geometric structure but demonstrate less alignment with the original stress directions. This discrepancy arises because the original stress field is smoothed, which alters the stress field globally and particularly eliminates unwanted degeneracies.
@@
The visualization of the 3D conforming lattice structure reveals several stitching edges in the interior that do not align with any of the dominant stress directions. Additionally, the number of longer edge sequences that consistently follow one of the PSLs is lower compared to the PSL-guided infill and the volumetric Michell trusses. This behavior might be attributed to the approach allocating too much material to maintain a consistent edge graph rather than prioritizing stiffness. Oscillating load paths, represented by inconsistent edge sequences, may lead to less effective load transitions and, consequently, reduced stiffness.
4.5 Variable Load Structural Analysis
To shed light on how well a design optimized for a specific load case can resist forces applied from a different direction, SGLDBench is used in the following way: First, all six infill design methods are applied with the same boundary condition to compute six different designs, and the von Mises stress field in the optimized material field is computed. Then, the applied forces are changed, and the von Mises stress field is recomputed for all designs. The designs, color-coded with the von Mises stress under the varied forces, are shown in Fig. 13.
Density-based TO consistently shows higher compliance under the new force conditions. While the compliance increases only slightly with a slight change of the force direction, larger directional changes result in a significant loss in stiffness. The coloring with the von Mises stress shows where the critical structures occur under the new forces and have the potential to break. A significant re-distribution of the von Mises stresses indicates a significant loss in stiffness.
All other designs show the same slight change in stiffness as porous infill when the force direction changes only slightly. For more significant changes in the force directions, however, all other designs perform significantly better than the one optimized with density-based TO. Interestingly, even the PSL-guided material layout, which is specifically aligned with the major stress directions occurring with the initial force setting, can significantly better resist the new force directions.
Designs that show a more uniform distribution of material throughout the domain, such as the Voronoi infill, the conforming lattice structures, and the volumetric Michell trusses, are much better at resisting varying forces compared to topology-optimized designs. When the force directions are varied significantly from their initial values, in some cases the stiffness becomes even higher. These are situations where the new forces are along some of the principal stress directions that have occurred in the initial stress field, and along which some of the structures have grown, so that the structure is bearing more normal stresses than shear stresses.
Figure 12: A clip plane reveals the interior structure of different designs. The closeup view shows the PSLs in the selected region, with the major, medium, and minor PSLs, respectively, shown in ocher, green, and blue.
@@ 4.6 Reproducing Stresses in the Solid Domain
Using Part shown in Fig. 13, SGLDBench is applied to analyze how well different designs reproduce the stress directions in the initial solid configuration. This analysis allows users to examine, for instance, whether force transmission through tension and compression occurs along the same load paths as in the initial solid. Since an infill design can, in principle, transmit forces along the most efficient load paths in the solid design or diverge to alternative paths, understanding the relationship between compliance and the reproduction of stress directions can provide valuable insights for improving infill designs.
Figure 13: Variable load structural analysis. Top row: All infills are optimized for the same boundary conditions (see inset with downward forces). Rows 2–4: Varying forces (Euler angles $\theta_{x},\theta_{y},\theta_{z}$ indicate angular deviation from the downward direction) are applied to the optimized infills. All designs are shown with their compliance $c$ and are color-coded according to the von Mises stress.
@@
Fig. 14 shows visualizations of the alignment fields computed for each infill. Iso-surface rendering with the color transfer function described in Sec. 3.4 is used to emphasize regions with high deviation in brown. TO produces an infill that, in many regions, aligns well with the major stress directions in the solid under load. However, significant deviations are observed in certain areas, particularly in thin structures and regions near the fixed elements where stress concentrations are highest.
Similarly, porous infill exhibits more pronounced deviations, which are distributed throughout the entire domain. This behavior arises from its space-filling material distribution, which emphasizes uniformity rather than alignment with the stress directions.
The stress deviations are highest in the Voronoi infill, as the initial stress directions are not considered in its edge-graph layout. Only when edges align with boundary elements, which are incorporated into the stress analysis of both the solid object and the infill, stress deviations are lower.
The material in the PSL-guided infill is distributed along the principal stress directions, resulting in good alignment with the initial stress field. However, as shown in Fig. 14, alignment accuracy decreases at the boundaries of the voxelized edges. This reduction in accuracy stems from discretization-induced inaccuracies in the stress simulation—a limitation shared by all graph-based approaches when voxelized structures are used for compliance analysis.
Additionally, all graph-based infills contain numerous edges aligned with the two principal stress directions not corresponding to the direction of the maximum absolute stress value. This can lead to high directional deviations along these edges.
@@
The deviation maps of the conforming lattice structure and the Michell trusses are very similar, showing the same low deviation patterns in the boundary region as the Voronoi infill. In both infills, there is still a considerable number of voxels with significantly different dominant stress directions than the solid object.
Figure 14: Stress-to-stress alignment. High (white) to low (brown) alignment between stresses in the solid and the optimized infill with respect to the same boundary conditions is shown. A view plane aligned clip plane reveals interior parts.